OpenStudy (anonymous):

Find a vector parallel to the line of intersection of the two planes x+2y-3z=7 and 3x-y+z=0.

OpenStudy (anonymous):

just a quick question are you doing matrices right now or it more of trig?

OpenStudy (anonymous):

it's for calc 3 we haven't done anything with matrices

OpenStudy (anonymous):

i tried to do an odd problem that was similar however i am just not getting it

OpenStudy (anonymous):

oh snap i was thinking about algebra lol kk

OpenStudy (turingtest):

I haven't seen this problem before, but I'm thinking that crossing the normal vectors of the planes should give a vector parallel to the intersection line does that make sense to you?

OpenStudy (turingtest):

\[\vec n_1\times\vec n_2=\left|\begin{matrix}\hat i&\hat j&\hat k\\1&2&-3\\3&-1&1\end{matrix}\right|\]

OpenStudy (turingtest):



OpenStudy (turingtest):

that's how I see it at least...